OpenStudy (anonymous):

PLZ PLZ HELP! MEDAL!! Which graph represents the solution to the given system? -4x + 2y = -11 and -2x + 2y = -15

OpenStudy (anonymous):

Graph 1

OpenStudy (anonymous):

Graph 2

OpenStudy (anonymous):

Graph 3

OpenStudy (anonymous):

And graph 4

OpenStudy (anonymous):

It would be the 4th graph because all you have to do is solve the two equations and put them in your calculator to see the graph. y=2x-5.5 y=x-7.5

OpenStudy (anonymous):

Thank you so much!! @IloveCharlie Anyone else?

OpenStudy (anonymous):

If you need to show work: -4x+2y=-11 you would add 4 onto the right side, and the whole equation by 2 2y=4x-11/2 so then you would get: y=2x-5.5 2x+2y=-15 2y=2x-15/2 y=x-7.5

OpenStudy (anonymous):

Thank you soooo much!!!!

OpenStudy (anonymous):

I have one more if anyone can help me!!

OpenStudy (anonymous):

Which graph represents this sytem of inequalities? y ≤ -3x - 2 y ≥ 2x + 1

OpenStudy (anonymous):

Graph 1

OpenStudy (anonymous):

Graph 2

OpenStudy (anonymous):

Graph 3

OpenStudy (anonymous):

Graph 4

OpenStudy (anonymous):

@him1618 @annas @cwrw238

OpenStudy (anonymous):

graph 2 is the answer because solutionset is (-2, -19/2)

OpenStudy (anonymous):

Thank you so so much!!

OpenStudy (anonymous):

this was answer to first

OpenStudy (anonymous):

Oh. So the other girl got the first question wrong>

OpenStudy (anonymous):

Oh yes first one is quiet simple you just have to get solution set for it to plot the graph & let me see the second one :)

OpenStudy (anonymous):

Oh okay! Thank you!!

OpenStudy (anonymous):

no problem, just a moment

OpenStudy (anonymous):

Alright!

OpenStudy (anonymous):

@annas You still there??

OpenStudy (anonymous):

answer is graph 1

OpenStudy (anonymous):

sorry was plotting graphs xD

OpenStudy (anonymous):

internet connection was down :/

OpenStudy (anonymous):

the correct answer for the first system would be graph 4 because the number before x is positive in both, therefore it must have a positive correlation.